OpenStudy (anonymous):

Solve the system algebraically: - x + y = 4 y = x^2 - 4x + 2

OpenStudy (unklerhaukus):

add x to both sides of the first equation, they you have two expressions for y which means they are equal equate them and solve

OpenStudy (anonymous):

x+4=x^2-4x+2 0=x^2-5x-2

OpenStudy (anonymous):

use the quadratic formula...